OpenStudy (anonymous):

solve this equation 7x^2+6=13

OpenStudy (anonymous):

x = -1 x = 1

OpenStudy (anonymous):

7x^2 +6 = 13 7x^2 = 13-6 7x^2 = 7 (7x^2)7 = 7/7 x^2 = 1 x = + or - 1